sin(theta) = cos(2theta + 12degrees)
find the acute angle that satisfies the equation

Answer by lwsshak3(11628)

sin(theta) = cos(2theta + 12degrees)
find the acute angle that satisfies the equation
***
consider cos and sin functions to be complementary:
theta+2theta+12˚=90˚
3theta=78˚
theta=26˚
..
Check:
sin(theta)=sin(26˚)=0.438...
cos(2 theta+12˚)=cos(64˚)=0.438...